Shift report 20010823 16:00-24:00

From: Apache (apache@pii3.brahms.bnl.gov)
Date: Thu Aug 23 2001 - 23:54:11 EDT

  • Next message: Ramiro Debbe: "Re: Shift report 20010823 16:00-24:00"

    ------------------------------------------------------------
    Date: Thu Aug 23 23:54:10 2001, Shift: 16:00-24:00
    Supervisor: JJG, with: EJ
    Log book#8 pages 10 - 14
    
    Shift Summary:
    ------------------------------------------------------------
    1. No beam today. No beam this evening, but plenty 
       of pedestals.
    2. Access to the IR during the day. 
    3. Ramiro operated on section 2 
       of T3 and surgically removed all life support systems 
       of the V plane. This means that T2 anode can be raised 
       to operating voltage (read also in the logbook).
    4. Dana and EB have completed the finger counters for the 
       FFS. They are now positioned just behind C1. The 
       counters are separated by 2.5". JJG and Erik have 
       cabled up , adjusted disc., applied HV and checked 
       that pulses come out to the FEE. See 
       details in logbook 8 and C1 logbook.
    5. HV has been set using the FFS1440 HV  pgrm to -1600V. 
       Need to include this in BRAVO (?), implement the data
       readout (FV?) and enable display in supermon(JIJ, KH?).
    6. FV and Betty + more have worked hard on the HPSS link. 
       FV says it is now operational. Since I have nothing to 
       transfer, I cannot verify this.
    7. It is now 11.25PM and RHIC says: heavy ions have been 
       restored to W-dump. This sounds ominous, but is perhaps
       a good sign. I will soon also retire to the B-dump.
    8. In any case, JJG leaves BNL tomorrow. Event though the  
       machine hasn't performed brilliantly (nice euphemism)
       the experiment is in pretty good shape, thanks to
       the many efforts of dedicated collaborators. The 
       data analysis is also coming well along, so things 
       look faily bright to me.
       Keep up the good work and lets all look forward to some
       exciting physics results. 
       Here ends the pastoral missive...Goodnight.
    ------------------------------------------------------------
    



    This archive was generated by hypermail 2b30 : Thu Aug 23 2001 - 23:54:44 EDT